# Training-Free Adaptive Quantization for Variable Rate Image Coding for Machines

### v2

@@ Abstract
Image Coding for Machines (ICM) has become increasingly important with the rapid integration of computer vision technology into real-world applications. However, most neural network-based ICM frameworks operate at a fixed rate, thus requiring individual training for each target bitrate. This limitation may restrict their practical usage. Existing variable rate image compression approaches mitigate this issue but often rely on additional training, which increases computational costs and complicates deployment. Moreover, variable rate control has not been thoroughly explored for ICM. To address these challenges, we propose a training-free quantization strength control scheme that enables flexible bitrate adjustment. By exploiting the scale parameter predicted by the hyperprior network, the proposed method adaptively modulates quantization step sizes across both channel and spatial dimensions. This allows the model to preserve semantically important regions while coarsely quantizing less critical areas. Our architectural design further enables continuous bitrate control through a single parameter. Experimental results demonstrate the effectiveness of our proposed method, achieving up to 11.07% BD-rate savings over the non-adaptive variable rate baseline.
@@ II-C Variable Rate Image Compression
While conventional LIC models are trained for a fixed rate, some studies have explored variable bitrate control to accommodate diverse bandwidth and quality requirements. Early works adopt Recurrent Neural Network (RNN)–based frameworks [20][22], though they often require high computational cost due to their progressive coding. Subsequent studies introduce conditional autoencoder–based approaches [23][25], followed by quantization-based rate control methods [26], [27]. For instance, QVRF [26] employs a quantization regulator to manage the overall quantization error of latent representations, enabling both discrete and continuous bitrate adjustment. Another line of work explores mask-based selective coding. J. Lee et al. [28] achieve variable rate compression by selectively encoding essential latent representations using a learned 3D importance map generated from the hyperprior outputs.
@@ III-A Preliminary Findings
We also examine the spatial characteristics of the scale parameters, as illustrated in Fig. 2. Larger scale values are observed along object boundaries, while smaller values appear in other regions such as object interiors and backgrounds. Given that SA-ICM is designed to preserve edge information effectively, these results indicate that the scale parameter reflects the region-wise importance relevant to recognition tasks. This observation forms the key intuition behind our proposed method.

### v3

@@ Abstract
Image Coding for Machines (ICM) has become increasingly important with the rapid integration of computer vision technology into real-world applications. However, most neural network-based ICM frameworks operate at a fixed rate, thus requiring individual training for each target bitrate. This limitation may restrict their practical usage. Existing variable rate image compression approaches mitigate this issue but often rely on additional training, which increases computational costs and complicates deployment. Moreover, variable rate control has not been thoroughly explored for ICM. To address these challenges, we propose a training-free framework for quantization strength control which enables flexible bitrate adjustment. By exploiting the scale parameter predicted by the hyperprior network, the proposed method adaptively modulates quantization step sizes across both channel and spatial dimensions. This allows the model to preserve semantically important regions while coarsely quantizing less critical areas. Our architectural design further enables continuous bitrate control through a single parameter. Experimental results demonstrate the effectiveness of our proposed method, achieving up to 11.07% BD-rate savings over the non-adaptive variable rate baseline. The code is available at https://github.com/qwert-top/AQVR-ICM.
@@ II-C Variable Rate Image Compression
While conventional LIC models are trained for a fixed rate, some studies have explored variable rate methods to accommodate diverse bandwidth and quality requirements. Early works adopt Recurrent Neural Network (RNN)–based frameworks [20][22], though they often require high computational cost due to their progressive coding. Subsequent studies introduce conditional autoencoder–based approaches [23][25], followed by quantization-based rate control methods [26], [27]. For instance, QVRF [26] employs a quantization regulator to manage the overall quantization error of latent representations, enabling both discrete and continuous bitrate adjustment. Another line of work explores mask-based selective coding. J. Lee et al. [28] achieve variable rate compression by selectively encoding essential latent representations using a learned 3D importance map generated from the hyperprior outputs.
@@ III-A Preliminary Findings
We also examine the spatial characteristics of the scale parameters, as illustrated in Fig. 2. Larger scale values align with object boundaries, while smaller values appear in other regions such as object interiors and backgrounds. Given that SA-ICM is designed to preserve edge information effectively, these results indicate that the scale parameter reflects the region-wise importance for recognition. This observation forms the key intuition behind our proposed method.